Please help me solve this age problem:
The sum of Sonny and Cely's ages is 81. 3 years ago Sonny's age is Cely's age now. How old will they be 2 years from now?

Answer by benbjlin(6)

Please help me solve this age problem:
The sum of Sonny and Cely's ages is 81. 3 years ago Sonny's age is Cely's age now. How old will they be 2 years from now?
Sonny's Age=S. Cely's Age=C.
S+C=81
S-3=C
2 variables, 2 equations, solve:
1. Substitute C in S+C=81 becomes: S+(S-3)=81
2. Simplify: 2S-3=81
3. Add 3 to both sides: 2S-3+3=81+3 becomes 2S=84
4. Divide by 2: 2S/2=84/2 becomes S=42
5. Solve for C by using S+C=81: 42+C=81 becomes C=39.
6. Question asks how old they will be 2 years from now. So answer is S+2 and C+2: Sonny in 2 years = 44 years, Cely in 2 year = 41 years.